A carpenter caut an 8-foot board into two pieces. The length of one of the pieces must be three times the length of the other piece.
Make a drawing and write a segment addition equation to model the board lengths.

Answer by checkley77(12569)

x+(3x)=8
4x=8
x=8/4
x=2 ft. length of the shorter cut.
8-2=6 ft. for the longer cut.
Proof:
2+6=8
8=8